### v2

Abstract
If-then rules are widely used to explain machine learning models; e.g., “if employed = no, then loan application = rejected.” We present the first proposal to apply rules to explain the emerging class of large language models (LLMs) with retrieval-augmented generation (RAG). Since RAG enables LLM systems to incorporate retrieved information sources at inference time, rules linking the presence or absence of sources can explain output provenance; e.g., “if a Times Higher Education ranking article is retrieved, then the LLM ranks Oxford first.” To generate such rules, a brute force approach would probe the LLM with all source combinations and check if the presence or absence of any sources leads to the same output. We propose optimizations to speed up rule generation, inspired by Apriori-like pruning from frequent itemset mining but redefined within the scope of our novel problem. We conclude with qualitative and quantitative experiments demonstrating our solutions’ value and efficiency.
keywords:
Explainability, Rule-Based Explanations, Large Language Models, Retrieval-Augmented Generation
1 Introduction
Advances in large language models (LLMs) have yielded increasingly sophisticated artificial intelligence (AI) systems capable of accomplishing a wide variety of tasks. These advances are enabled not only by the increasing size and architectural complexity of the models, but also by innovations in how LLMs incorporate knowledge. One such prominent innovation is retrieval-augmented generation (RAG), by which the static, pre-trained knowledge of an LLM is augmented with dynamic, arbitrary knowledge retrieved at inference time. Although RAG enables more control over an LLM’s data, the provenance of information in the LLM’s response is further complicated as a result.
The emerging field of explainable artificial intelligence (XAI) has been addressing issues related to understanding the behavior of complex models, including LLMs. Many XAI methods focus on the simpler problem of local explanation, which attempts to explain specific model predictions rather than general model behavior. One popular local XAI methodology is feature attribution, which explains a model output by scoring the relative importance of the input features. For RAG systems, context attribution can assign scores to each piece of information retrieved by an LLM system at inference time [contextcite]. However, feature attribution is not actionable (e.g., to achieve recourse).
Counterfactual explanations address this problem, explaining a model’s output by citing a minimal input perturbation that would induce an alternative output [wachter2017counterfactual]. In RAG, a counterfactual explanation might reveal that the LLM would have generated a different answer if some piece of information were not provided or retrieved at runtime [RAGE]. However, there may be many minimal counterfactuals for a given prediction, which may overwhelm end users. Explanations using if-then rules mitigate these problems by summarizing patterns observed across a model’s predictions [anchors, MACHA2022101209, lore], but have not yet been considered in the context of LLMs with RAG.
We fill this research gap and present a rule framework for explaining the output of black-box RAG systems using rules that formulate input predicates based on the presence of RAG sources provided to the LLM. Let us illustrate using a healthcare example. Suppose that a clinic is using a RAG system connected to a medical database, and a medical research assistant working at the clinic asks the LLM, “What treatment is most effective for Long COVID fatigue?” Here, a rule could reveal that “Whenever ‘Document 123’ is included in the RAG sources, the LLM always recommends an unsafe Long COVID treatment.” (A rule’s antecedent could likewise be when a given document is omitted from the RAG sources.) This rule helps curators of the medical database flag documents such as Document 123 that, unbeknownst to them, advocate for unsafe treatments. In this case, Document 123 may advocate for a treatment that was thought to be effective until recently, while newer studies have shown it is both ineffective and unsafe for certain demographics. Our framework empowers developers, auditors, policy makers, and other end users to seek rules for any identifiable output condition, such as responses that are incorrect, contain misinformation, or reflect negative sentiment [han2024medical].
Our main contributions are as follows.
  1. 1.
    We propose a rule-based explanation formulation for RAG, the first of its kind. We offer two variants that predicate on the basis of retaining or omitting features (RAG sources). A rule explains the output returned by an LLM by identifying conditions that consistently hold in practice whenever certain features are retained or omitted.
  2. 2.
    We design efficient search algorithms to mine rule-based explanations, exploiting intuitions of rule validity to prune the search space and enable early termination of the search.
  3. 3.
    We conduct an experimental evaluation of our explanation search algorithms, presenting case studies showing the value of rule-based explanations of RAG systems for question answering, and highlighting the efficiency improvements due to our pruning techniques.
The rest of the paper is structured as follows: Section 2 reviews related work; Section 3 introduces our rule formulation; Section 4 illustrates our rule mining algorithms; Section 5 presents an experimental evaluation; and Section 6 concludes with future research directions.
2 Related Work
In the XAI literature, many methods have been proposed to explain individual model predictions, collectively known as local explanations. Many local methods adopt an approach known as feature attribution, exemplified by methods such as LIME [ribeiro2016should] and SHAP [shap], which explain a model by scoring input features according to their relative influence on the model’s output. Though intuitive, these explanations are not directly actionable, and have been shown to be potentially unreliable, misleading, and manipulable [Kindermans2019, advlime:aies20, shap-edu-problems]. In contrast, counterfactual explanations explain a model by perturbing the input features in a minimal way to change the original output [wachter2017counterfactual, verma2024counterfactual]; e.g., a declined loan application would have been approved if the applicant’s income were 20 percent higher. Counterfactual perturbations aim to produce actionable insights that facilitate recourse, but many valid counterfactuals may exist, potentially contradicting each other and overwhelming users. Some methods therefore seek to sample diverse counterfactual perturbations or prioritize those deemed more feasible [mothilal2020dice, NEURIPS2020_c2ba1bc5].
At the cost of increased computational complexity, rules mitigate the above problems by summarizing patterns observed across a model’s predictions [anchors, MACHA2022101209, lore, rudinrules]; e.g., “if employed = no, then loan application = rejected,” regardless of the values of the other features. Recent works have established that rules essentially subsume counterfactuals, forming a duality, since both make assertions over a model’s input features [geng2022computing, guidotti2024stable]. Rule-based explanations have been defined as conjunctions of predicates over the features of an input, for which any matching input always leads to the model producing the same output [rudinrules]. Several methods, such as Anchors [anchors], have applied similar definitions for the purposes of XAI. In contrast to our work, their rules are approximate (i.e., hold with high confidence), not exact, and do not provide an interface to instantiate their rules’ input predicates for modern LLMs or RAG systems.
These rule formulations build implicitly upon association rule formulations established in the rule mining literature [rules-1993, apriori, most-interesting-rules]. These formulations gave rise to many influential rule mining algorithms that search for frequent itemsets and valid association rules, most notably the Apriori algorithm [apriori] and more-efficient successors like Eclat [eclat] and FP-growth [fpgrowth]. The rule mining methods we present in this work take algorithmic inspiration from Apriori, but have several key modifications for a fundamentally different problem: explaining AI model predictions. While Apriori iterates over observed itemsets, we instead iterate over the powerset of possible model inputs (features), eliminating any concept of support (see Definition 2) in rule validity. This enables us to seek sets that satisfy arbitrary predicates over model outputs, as opposed to those observed with some minimum frequency, as in the original Apriori algorithm. Our algorithms also leverage the concept of monotonicity (i.e., the Apriori property) to prune the search space, but we propose a new definition that operates on model output predicates rather than itemset support.
Naturally, many recent XAI works have focused on explaining LLMs, RAG, and other related phenomena, adapting general XAI methodologies as needed. Global explanation disciplines such as mechanistic interpretability attempt to identify low-level circuits that form among LLM neurons [autocircuits, NEURIPS2024_a8f7d43a, NEURIPS2024_ad217e0c]. However, these insights are intended primarily for scientific (i.e., expert) understanding. One intuitive explanation format for LLMs and RAG systems are citations, which can be generated at inference time using specialized training and fine-tuning methods [sourceaware2024, aly-etal-2024-learning] or post-hoc using various citation recommendation strategies [farber2020citation, maheshwari-etal-2025-citefix, ceg]. However, these methods are subject to the critical limitation of hallucination [agrawal-etal-2024-language], as their predictive approach to generating citations can be highly inaccurate, even with state-of-the-art models [citeme, cao-wang-2024-verifiable, byun-etal-2024-reference, day2023preliminary, alce]. Meanwhile, recent works have adapted feature attribution [contextcite] and counterfactual methods [RAGE] to explain source importance in RAG, assigning importance weights to the retrieved sources or generating counterfactuals by removing some sources to change the LLM’s output, respectively. In contrast to our rule-based approach, these recent methods are prone to the same limitations as their general methodologies, and may fail to produce concise and actionable insights that are robust and guaranteed.
3 Rule-Based Explanation Formulation
In this section, we introduce our novel rule-based explanation formulation. We start by reviewing the design space. We then provide a mathematical formalization of our rule-based explanations, clarifying how the hierarchical nature of our rules’ validity leads to complementary variants that either retain or omit inputs. Finally, we explain how rule validity can be defined recursively, and interpreted using a lattice data structure.
3.1 Rule Design
In the data mining literature, a rule expresses a conditional “if-then” relationship where satisfying the “if” part (the antecedent) implies that the “then” part (the consequent) is also satisfied [most-interesting-rules]. One intuitive adaptation for rules describing a prediction $y=M(x)$ might define the antecedent in terms of the input $x$ to the model $M$, and the consequent in terms of the resulting output $y$. This would yield rules of the form “if $x$ then yy”. We adopt this general orientation due to its clear alignment with the primary goal of local XAI: understanding the effect of model inputs on model outputs. This orientation has also been adopted in recent XAI works, wherein rules explain model predictions by uncovering significant patterns in said predictions. In this XAI setting, rules are better known as rule-based explanations.
Definition 1 (Rule-Based Explanation).
A rule-based explanation is a conjunction of predicates over the features of some model $M$, where for any input $x^{*}$ satisfying all predicates, the model always generates some target output $y^{*}$ [geng2022computing, rudinrules].
In the XAI setting, the antecedent is an assertion over a conjunction of input predicates, and the consequent is an assertion over a model output. In the literature, rules are often expressed and mined in terms of some set of observed instances, which would correspond to a model’s predictions in our setting. Given a set of observed instances, a rule may be judged by the number of observed instances that match both sides of the rule, or by the proportion of instances whose consequent is satisfied when the antecedent is satisfied. These empirical measurements, which are often used as a proxy for rule quality, are known formally as the support and confidence of a rule, respectively.
Definition 2 (Rule Support).
Given a set of observed instances, the support of a rule is the proportion of instances that satisfy the rule.
Definition 3 (Rule Confidence).
Given a set of observed instances, the confidence of a rule is the conditional probability that its consequent is satisfied, measured over all instances in the set that satisfy its antecedent.
In settings that observe such instances, rules are considered more informative when they have higher support (i.e., the rule is more likely to implicate an instance) and higher confidence (i.e., the rule is more likely to hold). In this work, we do not seek rules that characterize observed data, and thus do not presume a setting wherein a set of observed instances is available. Instead, we seek rules that characterize the full model input space, and therefore operate over the set of all possible antecedents. It follows that the classical notion of support does not apply in our setting, though it is nonetheless equal to the reciprocal of the number of possible antecedents. Furthermore, we require that all valid rules hold with certainty (i.e., confidence of 100%), but note that future work could build upon existing rule methods that guarantee minimum confidence levels [anchors].
In practice, ambiguity still remains when applying this formulation, as different models, features, problems, and user preferences necessitate different predicates. We now propose a concrete implementation for the RAG setting.
Antecedent: For our rules, the choice of antecedent input predicate(s) is critical, as it will determine their explanatory effectiveness and the computational complexity of mining them. Ideally, for maximum explanatory power, the antecedent of a rule should implicate as many inputs as possible (i.e., wide coverage). For example, the antecedent “If the LLM is given a set of sources that includes ‘Document 123’ ” covers more of the feature space than “If the LLM is given ‘Document 123’ ”. Likewise, the chosen antecedent predicate must also be compatible with and reasonable for the modality of the input features. While categorical features may be constrained to specific values (e.g., marital status is “single”), and numeric features to a range of values (e.g., credit score is $\geq 600$), the ideal definition for the textual inputs of language models is less clear.
Drawing inspiration from XAI literature, we contend that feature ablation (i.e., removing text segments) is a suitable perturbation strategy for explaining language models, and thus formalize antecedent predicates based on feature presence. In the XAI literature, ablation is a well-founded strategy for inferring feature importance [hameed:basedxai:2022], and has seen applications in works that explain language models [RAGE, contextcite] and models generally [shap]. We assume that the user partitions the input $x$ into two components $x=(\mathbf{s},\mathbf{c})$, where $\mathbf{s}$ represents some feature subset of interest $\mathbf{\mathbf{s}}=\{s_{1},s_{2},\dots,s_{n}\}$ derived from the input (e.g., RAG document sources). In our rule formulation, the antecedent will be defined in terms of the presence (or similarly, the absence) of inputs in $\mathbf{s}$, while the remaining inputs in $\mathbf{c}$ are ignored. By separating $\mathbf{s}$ and $\mathbf{c}$ (i.e., $y=M(\mathbf{s^{\prime}},\mathbf{c})$), we enable the user to focus their rules and explanations on small, interesting subsets of the input feature space, while reducing the complexity of our rule mining algorithms.
Consequent: We contend that the ideal choice of consequent is ultimately dictated by the user and their motivation for seeking rules. Therefore, the consequent of our rules asserts only that the output meets some target condition defined explicitly by the user. In essence, the target condition is an arbitrary output predicate that could test for any identifiable property (e.g., presence of misinformation, validity of in-text citations). We expect the user to define an output predicate function $O:\mathcal{Y}\rightarrow\{0,1\}$ that returns 1 when a model output $y\in\mathcal{Y}$ meets these conditions of interest. By parameterizing the consequent, we empower the user to seek rules that implicate not just the presence of a specific output (as is common in many XAI and rule-based works), but more sophisticated output characteristics (e.g., presence of hallucination, bias, sentiment).
To define an output predicate function $O$, the user must effectively implement a classifier that inspects an arbitrary LLM output and judges whether it meets the conditions of interest. A simple predicate might compare the answer for equality against some target answer, returning 1 if the LLM answer is equal to the target answer (and 0 otherwise). A more specific predicate might analyze the LLM answer for correctness; however, this would only be possible in settings where the correct answer is known upfront (e.g., benchmarking). For nuanced semantic assertions (e.g., judging citation validity), or those that must be robust to verbose or unstructured LLM outputs (e.g., spotting inappropriate content), a predicate function could adopt an “LLM-as-a-judge” implementation. Given detailed instructions and necessary context (e.g., ground truth information), an LLM can be a highly effective proxy for human judgment of LLM responses [chiang-lee-2023-large, judging-llm-judge]; however, stochasticity introduced during the decoding process (e.g., token sampling) can compromise reproducibility.
Figure 1: This diagram illustrates an example where the user asks a RAG system to recommend an effective treatment for Long COVID, given three retrieved sources. The rules generated by our algorithm (formalized in Section 4) indicate that retaining source $s_{2}$ consistently leads the LLM to recommend treatments known to be ineffective.
Example: Consider the following running example, illustrated in Figure 1. Suppose a user wants to debug incorrect predictions made by an LLM that is being used to suggest treatments for medical illnesses. The LLM is configured to use RAG, meaning that the LLM’s pre-trained knowledge is supplemented with external knowledge sources $\mathbf{s}$ (relevant medical documents) to help ground its answers. The user asks the LLM to recommend an effective treatment for Long COVID, and is investigating why an incorrect answer “calcium supplements” is consistently returned. Let $x$ represent the corresponding prompt submitted to the LLM $M$ by the user, and $y$ represent its incorrect response, as illustrated in Figure 1. All inputs in the prompt other than $\mathbf{s}$ (i.e., the question and LLM instructions) are captured in $\mathbf{c}$, and will not be part of any rules.
Suppose that the user wants to know whether any patterns emerge among the top-3 retrieved medical documents (sources) when ineffective treatments (i.e. incorrect answers) are observed. To make this intent explicit, the user defines a function $O_{incorrect}$ that returns 1 if $y$ recommends a treatment not known to be effective for the indicated condition, and 0 otherwise. The user implements $O_{incorrect}$ by constructing a prompt to be posed to a separate LLM. It contains a list of treatments known to be clinically effective for Long COVID, and asks it to return 1 if the observed response suggests a treatment not in this list. This definition is illustrated in Figure 1, which shows how the input $x$ and output predicate $O_{incorrect}$ (along with $M$ and $y$) are subsequently provided to our rule mining algorithm. The algorithm proceeds by considering all possible combinations of the three medical documents, performing inference for certain combinations (as necessary), then finally deriving patterns across combinations that satisfied $O_{incorrect}$.
As shown in Figure 1, one rule found to be valid under $O_{incorrect}$ asserts that “if $\mathbf{s}$ retains all sources in $\{s_{2}\}$, then $M$ returns an ineffective treatment.” Other valid rules are found, but they are simply more-specific versions of this rule (i.e., implicating supersets of $\{s_{2}\}$). If the user were instead interested in patterns that lead to a correct answer, they could define a reciprocal predicate $O_{correct}$, potentially producing a symmetric rule such as “if $\mathbf{s}$ omits all sources in $\{s_{2}\}$, then $M$ returns an effective treatment.” In either case, the rule suggests that the presence of $s_{2}$ is important for the LLM to recommend an ineffective treatment.
3.2 Rule Formulation
The concepts of presence and absence have a reciprocal relationship. Suppose that we have a subset of the input set, $\mathbf{s^{\prime}}\subseteq\mathbf{s}$, and that we aim to characterize the contents of $\mathbf{s^{\prime}}$ w.r.t. $\mathbf{s}$. Retaining all elements from $\mathbf{s}$ that appear in $\mathbf{s^{\prime}}$ is equivalent to omitting all elements from $\mathbf{s}$ that appear in the complement of $\mathbf{s^{\prime}}$, $(\mathbf{s}\setminus\mathbf{s^{\prime}})$. Conversely, omitting all elements from $\mathbf{s}$ that appear in $\mathbf{s^{\prime}}$ is equivalent to retaining all elements in $(\mathbf{s}\setminus\mathbf{s^{\prime}})$. This duality admits two variants of our rule-based explanations: retention and omission rules.
Definition 4 (Retention Rule).
A retention rule for some output predicate $O$ and implicated input subset $\mathbf{s^{\prime}}\subseteq\mathbf{s}$ guarantees that $\forall\mathbf{t}\in\mathcal{P}(\mathbf{s}),\ \mathbf{s^{\prime}}\subseteq\mathbf{t}\Rightarrow O(M(\mathbf{t},\mathbf{c}))=1$.
Definition 5 (Omission Rule).
An omission rule for some output predicate $O$ and implicated input subset $\mathbf{s^{\prime}}\subseteq\mathbf{s}$ guarantees that $\forall\mathbf{t}\in\mathcal{P}(\mathbf{s}),\ (\mathbf{t}\cap\mathbf{s^{\prime}}=\emptyset)\Rightarrow O(M(\mathbf{t},\mathbf{c}))=1$.
A retention rule identifies a subset of the input set that, when retained, always satisfies the output predicate. Similarly, an omission rules identifies a subset that, when omitted, always satisfies the output predicate. Retention and omission rules will often correlate with outputs that satisfy reciprocal (i.e., complementary) output predicates. Therefore, users will often use two separate, reciprocal output predicates when explaining a single input (e.g., $O_{correct}$ and $O_{incorrect}$). Generally, both rule types identify critical inputs in $\mathbf{s}$ whose presence is necessary to ensure some condition holds.
Revisiting our RAG example from Figure 1, recall that the user can obtain rules of two symmetric types, either in terms of answer correctness or incorrectness. In this example, the retention rule formulation aligns with $O_{incorrect}$, and the omission rule formulation aligns with $O_{correct}$. That is, the aforementioned retention rule under $O_{incorrect}$ asserts that, across all combinations of sources in $\mathbf{s}$, the LLM’s answer is always incorrect when the combination retains all sources in $\mathbf{s^{\prime}}=\{s_{2}\}$. Likewise, the complementary omission rule under $O_{correct}$ would assert that the answer is correct anytime the combination omits $\mathbf{s^{\prime}}=\{s_{2}\}$.
To guarantee these assertions for a subset $\mathbf{s^{\prime}}\subseteq\mathbf{s}$, one must verify that the output predicate remains satisfied for all possible subsets of $\mathbf{s}$ wherein the inputs in $\mathbf{s^{\prime}}$ are retained (or omitted, respectively). Following the sources $\mathbf{s}$ defined in Figure 1, where $|\mathbf{s}|=3$, validating either $\{s_{2}\}$ rule would require evaluating the LLM output corresponding to each subset in $T=\{\{s_{2}\},\{s_{1},s_{2}\},\{s_{2},s_{3}\},\{s_{1},s_{2},s_{3}\}\}$. To validate the $\{s_{2}\}$ retention rule against $O_{incorrect}$, one would need to ensure that $\forall\mathbf{t}\in T$, $O_{incorrect}(M(\mathbf{t},\mathbf{c}))=1$. The “Rule Miner” module in Figure 1 depicts these as blue nodes with a solid border, which indicates their validity. To validate the $\{s_{2}\}$ omission rule against the applicable predicate $O_{correct}$, one must ensure that $\forall\mathbf{t}\in T$, $O_{correct}(M(\mathbf{s}\setminus\mathbf{t},\mathbf{c}))=1$. In the next subsection, we formalize a definition of rule validity using recursion, which informs the design of our novel rule search algorithms.
3.3 Rule Validity
Under our definitions, the validity of a rule can be defined recursively. A retention rule implicating $\mathbf{s^{\prime}}$ is valid iff the predicate $O$ is satisfied for all supersets of $\mathbf{s^{\prime}}$. Likewise, for an omission rule $\mathbf{s^{\prime}}$, $O$ must be satisfied for all subsets of its complement $(\mathbf{s}\setminus\mathbf{s^{\prime}})$. Without loss of generality, the validity of a rule requires output predicate satisfaction for all ancestors of $\mathbf{s^{\prime}}$, including $\mathbf{s^{\prime}}$ itself. The set of ancestors can be defined recursively by repeatedly applying the parent relation, starting from $\mathbf{s^{\prime}}$ and continuing until no further parents exist. Thus, valid rules have an intrinsically recursive relationship, as their validity depends on the total validity of all ancestors.
Definition 6 (Rule Validity).
Let $P_{\mathbf{s^{\prime}}}$ denote the set of parents of $\mathbf{s^{\prime}}$, i.e., the immediate supersets of $\mathbf{s^{\prime}}$ (for a retention rule) or the immediate subsets of $(\mathbf{s}\setminus\mathbf{s^{\prime}})$ (for an omission rule). We can define rule validity as a recursive predicate $V(\mathbf{s^{\prime}})=O(\mathbf{s^{\prime}})\wedge\forall\mathbf{t}\in P_{\mathbf{s^{\prime}}},V(\mathbf{t})$.
By definition, all ancestors must be validated before validating a descendant. A rule mining algorithm must first validate those rules that retain or omit the most inputs (i.e., fully recurse) before unwinding the call stack to consider rules that retain or omit fewer inputs. Since a rule implicates all ancestor rules whose representative subset is a superset of its own, rules representing smaller subsets will implicate more rules than those representing larger, meaning that descendant rules have greater coverage than ancestor rules. Under this interpretation, descendant rules can be seen as subsuming ancestor rules as the latter have less coverage (implicate fewer feature combinations).
Definition 7 (Rule Subsumption).
A rule $r_{1}$ subsumes another rule $r_{2}$ iff every predicate in $r_{1}$ also appears in $r_{2}$.
Moreover, rules that retain or omit fewer inputs (i.e., descendants) are considered to have greater explanatory power, since by definition, their output predicate holds over more inputs (i.e., greater coverage). Rules that implicate smaller feature subsets $\mathbf{s^{\prime}}$ are considered to be more minimal.
Definition 8 (Rule Minimality).
A rule $r_{1}$ is minimal iff there exists no rule $r_{2}$ such that $r_{2}$ subsumes $r_{1}$ and $r_{2}\neq r_{1}$.
In this work, we seek minimal rules, but operate under the understanding that less-minimal rules must be validated first. We acknowledge an inherent trade-off between rule minimality and algorithmic efficiency, as finding more minimal (higher coverage) rules ultimately requires validating more rules. Our algorithms (introduced in Section 4) ultimately prioritize minimality, exhausting all rule candidates and guaranteeing 100% confidence. However, to eliminate redundancy, non-minimal rules can be discarded before being returned to the user.
Figure 2: A lattice representing an input set $\mathbf{s}$ where $|\mathbf{s}|=4$. The node coloring illustrates an example run for the top-down breadth-first search performed by Mono Rule Miner (introduced in Section 4.1), which discovers valid rules (blue with solid border), invalid rules (red with long dashes), and candidates that can be pruned (grey with short dashes).
The full hierarchy of a given rule is a subset of the power set $\mathcal{P}(\mathbf{s})$. The power set, which contains all candidate rules that implicate $\mathbf{s}$, forms a data structure known as a lattice [abiteboul1995foundations], illustrated in Figure 2. Each of the $2^{|\mathbf{s}|}$ subsets in $\mathcal{P}(\mathbf{s})$ is a rule candidate; each node in the lattice is a candidate, and each edge delineates a subset / superset relationship between two candidates. The $\mathcal{P}(\mathbf{s})$ lattice constitutes the full search space for any rule mining algorithm under our formulation, and it contains $2^{|\mathbf{s}|}$ candidate rules. In the following subsections, we present two rule mining algorithms that efficiently navigate this lattice in search of all valid rules.
4 Rule Mining Algorithms
In this section, we introduce two search algorithms [^1] Algorithm implementations are available at https://github.com/joelrorseth/rag-llm-rule-miner. capable of mining our rule-based explanations, each with different advantages. First, we introduce the Mono Rule Miner, which can mine either retention or omission rules in a single pass. Second, we introduce the Dual Rule Miner, which can mine both rule types simultaneously in a single pass. Mono is more efficient in settings where a single rule type is sufficient, and can exploit greater search space pruning. For settings where patterns of both types are sought and compared, Dual is often more efficient than executing two Mono passes.
4.1 Mono Rule Miner
As discussed in Section 3.2, users will request rule-based explanations after obtaining a model output $y=M(\mathbf{s},\mathbf{c})$. To define the desired rule, the user will specify whether they seek retention or omission rules, and will supply a definition for an output predicate $O$. The goal of our algorithms is to discover all valid rules that implicate subsets of $\mathbf{s}$. This necessitates a thorough (and potentially exhaustive) search through the lattice, and subsumes the simpler problem of validating a single candidate rule.
As exemplified in Figure 2, the lattice contains $|\mathbf{s}|+1$ levels that group rule candidates by their size (i.e., $|\mathbf{s^{\prime}}|$ for a rule based on $\mathbf{s^{\prime}}\subseteq\mathbf{s}$). Since rule validity is defined recursively w.r.t. all ancestors, we must generally navigate the lattice top-down to validate rules. Due to the recursive definition, validating any single rule can be accomplished using a depth-first search (DFS), starting from the rule in question and recursing to all its ancestors in the lattice. Since we seek all valid rules, simply executing DFS for each candidate rule (top-down) would incur significant repeated computation. Recomputation can be avoided by caching rule validity during a top-down traversal, reducing runtime complexity to $\mathcal{O}(2^{|\mathbf{s}|})$, but caching would impose an expensive $\mathcal{O}(2^{|\mathbf{s}|})$ space complexity.
Instead, we propose a top-down breadth-first search (BFS) implementation that finds all valid rules of a specific type in a single pass. We call this our Mono Rule Miner. The algorithm is detailed in Algorithm 1, and its design is informed by two key observations. The first is that our recursive definition of rule validity admits a memory-efficient dynamic programming solution. The second observation is that the recursive definition presents an opportunity for early termination when a rule is found to be invalid.
Dynamic Programming (Observation #1): To justify our dynamic programming solution, we observe that our recursive problem formulation is already a dynamic programming problem, since the validity of a rule is dependent on overlapping subproblems originating in its ancestor rules. Capitalizing on this, we can bound the number of cached subproblems (i.e., rule validity records) to a size that modestly improves over the naive $\mathcal{O}(2^{|\mathbf{s}|})$. Though the validity of a rule hinges also on the validity of all its ancestors, it is sufficient to hinge only on the validity of its immediate parents. This is true because the remaining ancestors must have already been validated when validating these parents. Therefore, we need only cache the validity of two levels worth of rules at any time (the current and previous level), resulting in a lower space complexity of $\mathcal{O}{|\mathbf{s}|\choose l^{*}}$ where $l^{*}=\lfloor|\mathbf{s}|/2\rfloor$ (the index of the largest level).
Early Termination (Observation #2): Our second observation arises from our definition of rule validity. Since a rule’s validity hinges on the validity of all ancestors, then if a rule candidate is found to be invalid, then none of its descendants can possibly be valid. This property is analogous to the downward closure property (Apriori property) from the data mining literature [apriori], and is key to “pruning” large subtrees of the lattice search space. Upon encountering an invalid rule, we can cache a record of its invalidity and avoid evaluating (i.e., performing model inference for) the child rules if and when we reach them during traversal. Furthermore, once a level has no valid rules, the entire search can terminate as all valid rules must have been found.
Algorithm 1 Mono Rule Miner
1:The input set $\mathbf{s}$
2:The remaining inputs $\mathbf{c}$
3:The output predicate $O$
4:The model $M$
5:The rule-based explanations
[…]
8:for $l^{\prime}\in[0,|\mathbf{s}|]$ do
9:  $l\leftarrow|\mathbf{s}|-l^{\prime}$ $\triangleright$ Compute current level size
10:  $S\leftarrow\{\mathbf{s^{\prime}}\subseteq\mathbf{s}\mid|\mathbf{s}^{\prime}|=l\}$ $\triangleright$ Get all current level nodes
11:  $Z\leftarrow\{\mathbf{z}\subset\mathbf{z^{\prime}}\mid\mathbf{z^{\prime}}\in Z\land|\mathbf{z}|=l\}$ $\triangleright$ Get children of invalidated parents
12:  $Q\leftarrow S\setminus Z$ $\triangleright$ Prune invalid nodes in current level
13:  for $\mathbf{s^{\prime}}\in Q$ do
14:   $y\leftarrow M(\mathbf{s^{\prime}},\mathbf{c})$ $\triangleright$ Invoke the model
15:   if $O(y)=1$ then
16:     $R\leftarrow R\cup\{\mathbf{s^{\prime}}\}$ $\triangleright$ Record valid rule
17:   else
18:     $Z\leftarrow Z\cup\{\mathbf{s^{\prime}}\}$ $\triangleright$ Record invalid node
19:   end if
20:  end for
21:end for
22:return $R$
Example: To illustrate how the the Mono Rule Miner works, we will now explain how Algorithm 1 navigates the lattice in Figure 2 to find all valid rules of a specific type. Once provided with inputs $\mathbf{s},\mathbf{c},O$ and $M$, the algorithm will initialize caches for valid and invalid rule candidates (nodes corresponding to subsets of $\mathbf{s}$), then proceed with a top-down level-order traversal (line 8). Supposing that $|\mathbf{s}|=4$ as illustrated in Figure 2, the traversal begins by gathering (line 10) and looping through (line 13) subsets of size 4 (i.e., any $\mathbf{s^{\prime}}\subseteq\mathbf{s}$ where $|\mathbf{s^{\prime}}|=4$, meaning only $\mathbf{s}$). Inside this loop, inference is performed on each of these subsets (line 14), and the predicate function $O$ is invoked with the output to determine rule validity (line 15). For this first subset $\{s_{1},s_{2},s_{3},s_{4}\}$, the output is found to satisfy $O$ (as indicated in Figure 2) and is added to the valid candidates cache.
The algorithm then proceeds to the next level containing subsets of size three. Since $\mathbf{s}$ is a valid rule and is the sole parent of all subsets on this level, then each subset may be a rule, and will be if it satisfies $O$. As reflected in Figure 2, all subsets are found to be valid rules except for $\{s_{1},s_{2},s_{3}\}$, whose descendants can now logically be ruled as invalid. When the algorithm discovers that this subset is invalid, only the subset itself is added to the invalid candidates cache (line 18), as descendants will be pruned dynamically upon navigating to subsequent levels. Proceeding to the next level with subsets of size 2, the invalid candidates cache (containing only $\{s_{1},s_{2},s_{3}\}$) is emptied and repopulated (lines 11 and 12) with the children of all members ($\{s_{1},s_{2}\}$, $\{s_{1},s_{3}\}$, $\{s_{2},s_{3}\}$), effectively pruning nodes from the subsequent evaluation loop.
Analysis: In the worst case, when all $2^{|\mathbf{s}|}$ rule candidates are valid, the algorithm must make $2^{|\mathbf{s}|}$ inference calls and use $\mathcal{O}({|\mathbf{s}|\choose l^{*}})$ space. However, in the opposite case, when no valid rules exist, our real-time pruning short-circuits the algorithm, requiring only a single inference call. Otherwise, the complexity varies depending on the number of valid rules, their location in the lattice, and many other indirect factors. These factors are analyzed empirically in Section 5.
4.2 Dual Rule Miner
As the name implies, the limitation of the Mono Rule Miner is that only rules of a single type are mined in its sole pass. In settings where comprehensive tests are demanded (e.g., LLM red teaming), obtaining rules for a pair of reciprocal rule types enables deeper understanding of the RAG sources and their influence. With rules of two reciprocal types, a user can compare whether a set of sources $\mathbf{s^{\prime}}$ was sufficient, both when retained and omitted, to induce an LLM output that meets a single target condition. If $\mathbf{s^{\prime}}$ is implicated in both types of rules, then users can be more confident in its absolute importance. However if $\mathbf{s^{\prime}}$ is implicated in only one rule type, users can infer that its influence is relative to the presence (or absence) of the other sources.
To find rules of both types using Mono, each would require a separate pass, and may utilize different output predicates. Despite potentially different predicates, the underlying BFS still navigates the same lattice, in the same order, for the same input set $\mathbf{s}$, regardless of the chosen type. Algorithmically, the key difference between the two types stems from the interpretation of the “current subset;” i.e., how it will be applied. For the current subset $\mathbf{s^{\prime}}\subseteq\mathbf{s}$, a retention rule must retain $\mathbf{s^{\prime}}$, but an omission rule must omit $\mathbf{s^{\prime}}$.
Guided by two additional observations (to follow), we devise a strategy to navigate the lattice in search of both types of rules, simply by interpreting each subset in opposite ways. We refer to the resulting algorithm as the Dual Rule Miner. It proceeds with the same top-down BFS lattice navigation as the Mono Rule Miner, finding rules of both types; however, some efficiency is sacrificed. Specifically, the benefits of our early termination (to reduce runtime complexity) and dynamic programming caching (to reduce space complexity) introduced in Section 4.1 are diminished.
Mask Representation (Observation #1): Recall from Section 3.2 that presence and absence have a reciprocal relationship. Omitting $\mathbf{s^{\prime}}$ is equivalent to retaining its complement w.r.t. $\mathcal{P}(\mathbf{s})$, ($\mathbf{s}\setminus\mathbf{s^{\prime}}$). While we can adopt this conceptually, we require an efficient representation to enable dual interpretation in a single algorithm pass. To solve this problem, we encode each lattice node (subset) as a mask, where each mask element reflects one of two possible states (True or False) for the input $s\in\mathbf{s^{\prime}}$ it represents. Masks permit a single canonical representation for each lattice node while simultaneously permitting two interpretations: one where True indicates retention of the corresponding input (and omission when False), and one where True indicates omission (and retention when False).
Pruning Adjustments (Observation #2): Although our Mono Rule Miner pruning strategy can still be used to prune invalid rules, we can only prune rules of the same type. Therefore, even if no further rules of one type can be found, lattice navigation will continue if rules of the other type may still be found. In short, Dual pruning opportunities can only be less common (and thus less effective overall) than Mono, since a subset must be invalid under both rule types before descendants can be pruned from the search. Moreover, once navigation reaches the bottom half of the lattice, the concrete subset for a given rule may have already been posed to the model $M$ when interpreted under the opposite type. In effect, this results in the $\mathcal{O}({2^{|\mathbf{s}|}})$ time complexity being multiplied by a constant factor of 2.
Algorithm 2 Dual Rule Miner
1:The input set $\mathbf{s}$
2:The remaining inputs $\mathbf{c}$
3:The retention output predicate $O_{ret}$
4:The omission output predicate $O_{omi}$
5:The model $M$
6:The rule-based explanations
[…]
11:for $l^{\prime}\in[0,|\mathbf{s}|]$ do
12:  $l\leftarrow|\mathbf{s}|-l^{\prime}$ $\triangleright$ Compute current level size
13:  $S\leftarrow\{\mathbf{s^{\prime}}\subseteq\mathbf{s}\mid|\mathbf{s}^{\prime}|=l\}$ $\triangleright$ Get all current level nodes
14:  $Z_{ret}\leftarrow\{\mathbf{z}\subset\mathbf{z^{\prime}}\mid\mathbf{z^{\prime}}\in Z_{ret}\land|\mathbf{z}|=l\}$ $\triangleright$ Get children of invalidated parents
15:  $Z_{omi}\leftarrow\{\mathbf{z}\subset\mathbf{z^{\prime}}\mid\mathbf{z^{\prime}}\in Z_{omi}\land|\mathbf{z}|=l\}$
16:  $Q^{\prime}\leftarrow\{(\mathbf{s^{\prime}},[\mathbf{s^{\prime}}\notin Z_{ret}],[\mathbf{s^{\prime}}\notin Z_{omi}])\mid\mathbf{s^{\prime}}\in S\}$ $\triangleright$ Pair nodes with their validity
17:  $Q\leftarrow Q^{\prime}\setminus\{(\mathbf{s^{\prime}},v_{ret},v_{omi})\in Q^{\prime}\mid\neg v_{ret}\land\neg v_{omi}\}$ $\triangleright$ Prune totally invalid nodes
18:  for $(\mathbf{s^{\prime}},v_{ret},v_{omi})\in Q$ do
19:   if $v_{ret}\land O_{ret}(M(\mathbf{s^{\prime}},\mathbf{c}))=1$ then $\triangleright$ Check for valid retention rule
20:     $R_{ret}\leftarrow R_{ret}\cup\{\mathbf{s^{\prime}}\}$
21:   else
22:     $Z_{ret}\leftarrow Z_{ret}\cup\{\mathbf{s^{\prime}}\}$
23:   end if
24:   if $v_{omi}\land O_{omi}(M(\mathbf{s}\setminus\mathbf{s^{\prime}},\mathbf{c}))=1$ then $\triangleright$ Check for valid omission rule
25:     $R_{omi}\leftarrow R_{omi}\cup\{\mathbf{s^{\prime}}\}$
26:   else
27:     $Z_{omi}\leftarrow Z_{omi}\cup\{\mathbf{s^{\prime}}\}$
28:   end if
29:  end for
30:end for
31:return ($R_{ret},R_{omi}$)
Example: The implementation of Dual Rule Miner is given by Algorithm 2. The algorithm requires the same inputs as Mono Rule Miner, except that two separate predicates $O_{ret}$ and $O_{omi}$ are expected for retention and omission rule types, respectively. Likewise, separate valid and invalid rule candidate caches are now maintained for each rule type; however, the familiar level-order traversal proceeds in a similar fashion (line 11). Each subset of the current level size is evaluated under both interpretations separately (lines 19 and 24), each applying their corresponding mask and being judged by their corresponding output predicate.
The main point where Dual diverges from Mono is the pruning step; i.e., just before looping through a level’s subsets. Since a subset must be invalidated under both types before it can be pruned, Dual must reinitialize the respective invalid candidate caches separately (lines 14 and 15), then prune only those subsets that do not appear in either (lines 16 and 17). For example, suppose that all subsets of size 4 and 3 have the same validity as indicated in Figure 2 regardless of type (i.e., $\{s_{1},s_{2},s_{3}\}$ is invalid for retention and omission). As the search continues to subsets of size 2, Dual prunes the same children as described in Section 4.1 for Mono, since these child subsets are invalid under both rule types.
Diverging from what is illustrated in Figure 2, suppose that $\{s_{1},s_{4}\}$ is found to be a valid retention rule, but an invalid omission rule. Consequently, each descendant of this subset may be a valid retention rule, but can never be a valid omission rule. The algorithm inserts $\{s_{1},s_{4}\}$ into the invalid omission candidates cache (line 27), resulting in the inclusion of its children $\{s_{1}\}$ and $\{s_{4}\}$ when reinitialized (line 15). Since this subset and its children were not included in both invalid caches, the evaluation loop will still “evaluate” these children. However, per-type validity is still tracked and associated with each evaluated subset (line 17), allowing the algorithm to “short circuit” and avoid unnecessary inference calls.
As with Mono Rule Miner, descendants of all invalid rule candidates will continue to propagate through the continuous reinitialization of their respective invalid candidate caches. As a result, once a descendant of a subset invalidated for one type is invalidated for the other type, all descendants of this now “totally invalidated” descendant will be pruned. Continuing our example, suppose that $\{s_{1}\}$ (already cached as an invalid omission rule via propagation) is found to be an invalid retention rule (via inference and judgment by $O_{ret}$). This subset will be added to the invalid retention rule candidates (line 22), and will therefore be present in both invalid caches, meaning that all descendants (only $\{\}$ in this case) will be pruned by the algorithm.
Analysis: The $\mathcal{O}({|\mathbf{s}|\choose l^{*}})$ space complexity of Mono Rule Miner (established in Section 4.1) remains largely intact here, but a constant multiple of 2 is again introduced to account for each rule type requiring its own rule validity cache. In case the user wishes to avoid the aforementioned repeated inference calls, Dual Rule Miner can cache model responses, indexing them by their canonical mask. Caching responses will raise the space complexity to $\mathcal{O}({2^{|\mathbf{s}|}})$. We allow the user to decide whether they would prefer repeated inference calls or higher memory use. In Section 5.2.2, we empirically study the proportion of overlapping inference calls in a real RAG setting, and analyze the effectiveness of Dual’s caching in different conditions.
5 Experimental Evaluation
In this section, we conduct an experimental evaluation [^2] The code to reproduce both quantitative experiments is available at https://github.com/joelrorseth/rag-llm-rule-miner. to analyze the efficiency and quality of our rule mining algorithms. We begin with a qualitative evaluation that describes a RAG case study in the domain of healthcare. Then, we present a quantitative evaluation studying the efficiency and scalability of our rule miners, applied in both real-world and synthetic settings. This is followed by another case study demonstrating a large-scale application of our rule miners in the domain of AI safety. Lastly, we present another quantitative evaluation studying the robustness of the rule miners and the rules they find across different commercial LLMs.
5.1 RAG Healthcare Case Study
Figure 3: A lattice corresponding to the case study presented in Section 5.1. Here, the minimal valid rule (lowest valid rule in the lattice, as per Definition 8) states that “if D2 and D4 are retained, then the LLM produces a response that is inconsistent with clinical evidence.”
We now present a case study that considers a real world scenario in the domain of healthcare. We document several XAI challenges that arise when using RAG-enabled LLMs in this setting, and analyze how our rule-based explanations help to rectify them. Our case study follows a technical user who is debugging a medical research system used by healthcare professionals (e.g., doctors, nurses) for the purposes of real-time research. One component of this system is a chatbot that enables end users to pose domain questions and obtain answers that are grounded in medical documents. This chatbot component is implemented using a retrieval-augmented LLM connected to a medical database.
The user is investigating a concerning trend that several doctors have identified: when posed with questions related to Long COVID, the chatbot often returns a misinformed answer that downplays the efficacy of pulmonary rehabilitation, despite clinical consensus of its benefits. The user suspects that one or more documents being retrieved and provided to the LLM contain medical misinformation, and that the LLM is ultimately grounding its responses in this misinformation. To dig deeper, the technical user wishes to understand how the LLM’s answer changes when seeded with different documents provided via RAG. Specifically, they wonder whether certain document combinations lead the LLM to produce responses inconsistent with the authoritative clinical evidence, and therefore seek rules that link specific documents to this inconsistency.
Suppose that the user is using a commercial LLM augmented with a prompt-based RAG pipeline, which works as follows. Upon receiving a question from the end user, the pipeline first prompts the LLM to generate one or more keyword queries that can be posed to a retrieval model to obtain documents that may prove helpful in answering the question. Then, for each generated keyword query, a proprietary retriever capable of scoring document relevance obtains the top-10 documents from the medical database (i.e., top-10 deemed most relevant to the query). The system establishes a final set of document sources $\left\{\text{D1},\text{D2},\text{D3},\text{D4},\text{D5}\right\}$ by sorting all retrieved documents by their relevance scores and keeping only the 5 highest scoring. These documents are inserted into a prompt that instructs the LLM to use them to answer the end user’s question.
The user decides to debug a specific question, “Is pulmonary rehabilitation effective for treating Long COVID symptoms?,” focusing first on the documents returned by the system. Upon review, the user confirms that the retriever is functioning correctly, and that the retrieved documents are relevant to the query. The user suspects that, despite their relevance, one or more of the retrieved documents may contain conflicting or inaccurate information that causes the LLM to produce a response inconsistent with clinical evidence. Therefore, the user seeks rules that will identify these problematic documents and link them to the inconsistent answers.
To generate the rules, the user defines an output predicate that uses a secondary LLM to judge the inconsistency of the response w.r.t. authoritative clinical evidence (supplied by the user). The user employs Mono Rule Miner to find all valid retention rules. As illustrated in Figure 3, several rules are returned, all originating from a single minimal rule that states the following: whenever documents $\left\{\text{D2},\text{D4}\right\}$ are retained, the LLM response is inconsistent. This suggests that, although neither D2 nor D4 are sufficient individually to induce an inconsistent answer (i.e., there is no valid rule for $\left\{\text{D2}\right\}$ or $\left\{\text{D4}\right\}$), their combined presence is.
With these new insights, the technical user can now flag these two documents for further review by a medical expert, and temporarily remove one or both from the database to prevent inconsistent answers. After the expert revises the documents to ensure the content is consistent with authoritative clinical evidence, the same rule-based explanations can aid in validating a successful redeployment. By running the rule miner again, the technical user should find that no rules implicating inconsistent answers exist, ensuring that the system is more likely to remain faithful to the clinical consensus.
5.2 Quantitative Evaluation: Rule Miner Efficiency
In this section, we present a quantitative experiment that evaluates the pruning improvements and scalability of our rule miners, measured over both real-world and synthetic data. We analyze rule miners’ efficiency both absolutely and relatively, discussing practical implications and directions for future improvement.
5.2.1 Experimental Setup
To test our rule miners in a modern setting, we design an experiment that tasks a RAG system with question answering (QA), seeking rules to explain its answer to specific questions. We focus specifically on a RAG QA scenario where a factual question is posed to an LLM along with a set of knowledge sources for reference. The LLM, $M$, is prompted to return the correct answer if it can be inferred from the knowledge sources, or “N/A” otherwise. The full implementation details of this “LLM-under-test” are provided in Appendix A. To realize this experimental setting, we utilize the HotpotQA dataset [hotpotqa], a QA benchmark for LLMs and other QA systems.
HotpotQA tests a model’s ability to answer multi-hop questions, which require consideration of multiple independent pieces of information (“hops”) to answer. HotpotQA questions have exactly two hops. An example multi-hop question with two hops might ask, “What is the chemical element named after the physicist who developed the theory of relativity?” Logically, answering this question requires one to first learn that Albert Einstein primarily developed the theory of relativity, then learn that the element named after him is Einsteinium. Each HotpotQA question is labeled with a single ground truth answer, a set of necessary sources required to infer the answer (constructed from HotpotQA “supporting facts”), and a set of distractor sources that are inconsequential to the determination of the answer. The sets of necessary and distractor sources are disjoint, and there are exactly $k=2$ necessary sources for each question (one for each hop).
As formalized in Section 3.2, our rule miners require a definition for the input set $\mathbf{s}$ and the output predicate(s) $O$. To decide on a formulation, we must motivate our search for rules. We adopt an intuitive problem definition wherein a user wishes to understand how the context sources provided to the LLM (via HotpotQA) affect the correctness of the LLM’s answer. Therefore, when posed with a specific HotpotQA question, we define $\mathbf{s}$ as the set of sources that we provide to the LLM. The question and other input components constitute $\mathbf{c}$, since they are not the focus of our present formulation.
Meanwhile, we define two complementary output predicates, $O_{consistent}$ and $O_{inconsistent}$, which attempt to classify answers predicted by the LLM based on their consistency w.r.t. the ground truth answer. The predicate function first attempts to judge a predicted answer’s correctness via deterministic string matching against the ground truth answer. If the answers do not match, they are posed to an LLM (gpt-5-mini-2025-08-07) that judges the equivalency of the candidate answer w.r.t. the ground truth answer. The full implementation details for the string matching and “LLM-as-a-judge” procedures are provided in Appendix A. Under this formulation, our algorithms will mine the following two types of rules:
  1. 1.
    Retention rules: Identify source subsets that, when retained within a RAG prompt, always induce an answer consistent with the ground truth answer.
  2. 2.
    Omission rules: Identify source subsets that, when omitted from a RAG prompt, always induce an answer inconsistent with the ground truth answer.
The goal of this quantitative evaluation is to measure the efficiency of our rule miners, which leverage novel pruning and early termination strategies, as compared to a naive method that does not. We propose that the efficiency of a rule miner can be quantified by the average proportion of lattice nodes that the rule miner actually visits. In this setting, “visiting” a node implies executing a single LLM inference call, therefore the number of calls executed by the rule miner can also quantify its efficiency. Expressing the number of nodes visited as a proportion allows these measurements to be averaged and consolidated over lattices of varying sizes (i.e., varying numbers of possible nodes).
Using these two metrics, we proceed as follows. We draw 50 QA pairs at random from the HotpotQA training set (hotpot_train_v1.1.json), which will ultimately be posed to a single “LLM-under-test” $M$ (gpt-5-mini-2025-08-07). For each question, we task each of our rule miners (Mono seeking retention rules, Mono seeking omission rules, and Dual seeking both) to mine all rules arising from posing said question to $M$. Given that $k=2$ (i.e., only two sources should be consequential to consistency), we anticipate more opportunity for pruning as additional distractor sources are considered and $|\mathbf{s}|$ increases. Therefore, we pose each question repeatedly to $M$ with increasingly large $\mathbf{s}$ (spanning $1\leq|\mathbf{s}|\leq 10$), exhausting necessary sources before drawing additional distractor sources.
To contrast the results of the real-world HotpotQA setting, we also compute results in a synthetic setting. In this setting, we control for several real-world factors that affect the true validity of rules found. We create a dataset containing mock sources (both necessary and distractor), then instantiate a mock LLM-as-a-judge and LLM-under-test $M$ that use them to predict consistent vs. inconsistent perfectly. In effect, this controls for the non-determinism and stochasticity of both LLMs, as well as for their propensity to make inaccurate and inconsistent predictions. We aggregate results over 1000 synthesized examples, and otherwise perform the same steps as in the HotpotQA setting. In the following, we discuss results and contrast them in the HotpotQA vs. synthetic settings.
5.2.2 Experimental Results: Pruning Improvements
(a) Rule miner lattice exploration as evaluated on a real LLM tasked with answering HotpotQA questions.
(b) Rule miner lattice exploration as evaluated in a synthetic setting that controls for stochasticity and consistency.
Figure 4: The average degree of pruning achieved by different rule miners, as measured by the average proportion of lattice nodes evaluated. These measurements are plotted over increasing lattice sizes, showing that our Mono and Dual rule miners prune more nodes as more distractor sources are considered (albeit at a slowing rate).
The results illustrated in Figure 4 demonstrate significant efficiency gains via pruning, avoiding many of the $2^{|\mathbf{s}|}$ nodes that a naive algorithm would evaluate. Generally, once all necessary sources are provided to the LLM (i.e., $|\mathbf{s}|\geq 2$), the rule miners deem additional sources inconsequential to the consistency of predicted answers and prune larger portions of the lattice. Figure 4(a) demonstrates that Mono Rule Miner can prune as much as 50% of the nodes when seeking retention rules in the HotpotQA setting, and as much as 70% when seeking omission rules. The synthetic setting results in Figure 4(b) reflect similar trends with slightly smaller rates, indicating clear asymptotic convergence at a 25% pruning rate for omission rules and 75% for retention rules. Across both settings, the proportion of this pruning increases with $|\mathbf{s}|$, albeit at a slowing rate.
Mono Rule Miner pruning rates vary significantly by rule type, but the trend flips in different settings: more omission rules are pruned than retention rules in the HotpotQA setting, but more retention rules are pruned in the synthetic setting. Naturally, the number of valid rules found mirrors this same pattern (see Figure 8 in Appendix B). In our RAG QA formulation, retention rules must identify source combinations sufficient to induce a consistent answer (e.g., combinations with both necessary sources), but omission rules need only identify a combination whose removal induces an inconsistent answer (e.g., any combination containing one or more necessary sources). Therefore, in a deterministic setting (like our synthetic setting), we expect more omission rules than retention, but the non-determinism introduced by using real LLMs with real data (like our HotpotQA setting) defies these expectations. Due to several factors discussed in Section 5.4.2, especially the fact that the LLM guesses consistent answers without all necessary sources, fewer omission rules hold than expected and more retention rules hold than expected.
(a) Mono vs. Dual Rule Miner lattice exploration as evaluated on a real LLM tasked with answering HotpotQA questions.
(b) Mono vs. Dual Rule Miner lattice exploration as evaluated in a synthetic setting that controls for stochasticity and consistency.
Figure 5: The average degree of pruning achieved by Dual Rule Miner as compared to two equivalent Mono passes (i.e., seeking retention rules, then omission). Plotted over increasing lattice sizes and w.r.t. 200% lattice exploration, Dual avoids reevaluating a significant proportion of lattice nodes, but to a lesser extent with additional distractor sources.
Figure 4 reports that, unsurprisingly, Dual Rule Miner evaluates most lattice nodes and evaluates more nodes than either Mono run individually. Importantly, it also confirms that Dual avoids evaluating any nodes more than once, a key advantage over the alternative strategy for finding both rule types: running Mono once for each rule type. Figure 5 provides a more accurate juxtaposition of pruning efficiency for Mono vs. Dual, rescaling “exploration” in terms of two lattice passes. Dual maintains an exploration rate near 100% in both HotpotQA and synthetic settings (decreasing and eventually converging at 75% in HotpotQA), while Mono evaluates nearly 200% (i.e., all nodes twice) in the worst case. As more distractor sources are considered, Mono’s exploration rate drops at a steady rate before eventually converging to Dual’s rate. In short, Dual is more efficient for smaller sets of sources.
5.2.3 Experimental Results: Scalability
Interpreting the same data under a different lens, Figure 6 visualizes and contrasts the scalability of each rule miner. Specifically, it plots the number of predictions made by each rule miner against the number of sources considered. Generally, all rule miners improve upon the naive algorithm that would evaluate all $2^{|\mathbf{s}|}$ lattice nodes. Moreover, the rate of scaling is consistent not only with the worst-case time complexity derived in Section 4, but is highly consistent across both the HotpotQA and synthetic settings.
By bounding this efficiency experiment to $1\leq|\mathbf{s}|\leq 10$, we also establish an upper bound on the number of sources reasonably supported by our rule miners at $|\mathbf{s}|=10$. We note that many real-world information retrieval (IR) systems (e.g., web search engines) seek the top-10 documents most relevant to users’ queries [JANSEN2006248, web-search-engines], making this an appropriate boundary for explaining real-world RAG LLMs that utilize such systems. Figure 6(a) shows that when $|\mathbf{s}|=10$ in the HotpotQA setting, Mono Rule Miner finds all omission and retention rules using approximately 250 and 500 predictions respectively, while Dual finds both using approximately 750. Therefore, when $|\mathbf{s}|=10$, the prediction count of Dual is similar to running Mono twice (minimal efficiency improvement), but as $|\mathbf{s}|$ decreases, Dual requires slightly less than double (moderate efficiency improvement). Given that many real-world RAG LLMs (e.g., LLMs tested in these experiments) can complete QA inference calls with sub-second response time, we assert that all rules can be found in a few minutes when $|\mathbf{s}|\leq 10$.
When $|\mathbf{s}|>10$, computing all rules can become intractable due to the exponential time complexity. This asymptotic boundary is well-studied in the association rule mining literature, and is largely unavoidable due to the exponential number of combinations that must be evaluated in the worst case (the entire lattice). The algorithms presented in this work adopt a design similar to the original Apriori algorithm [apriori], which ultimately navigates this exponential lattice level-order. More efficient successors to Apriori have emerged in the literature, and should inform successors of the present work, but the exponential time complexity remains. We point interested readers to works such as Eclat [eclat], whose depth-first traversal strategy improves I/O overhead, and FP-growth [fpgrowth], whose divide-and-conquer strategy avoids candidate generation.
In the next subsection, we present a case study that describes a method for tractably applying our rule miners when $|\mathbf{s}|>10$, so long as $k$ (the number of truly necessary sources) is a small constant (e.g., $k\leq 10)$. We contend that, although $k$ is not necessarily known upfront by users of RAG systems, it is unlikely that both $|\mathbf{s}|>10$ and $k>10$ in many practical settings. IR research has shown that in real-world search engines, the probability of a ranked document being relevant generally decreases monotonically as rank increases [analysis_rank_dist], suggesting that $k$ is likely small in RAG settings. Alternatively, rules can be mined for large $|\mathbf{s}|$ by partitioning them (i.e., find rules separately for smaller batches) or by merging them (i.e., find rules of coarser granularity).
(a) Rule miner predictions as evaluated on a real LLM tasked with answering HotpotQA questions.
(b) Rule miner predictions as evaluated in a synthetic setting that controls for stochasticity and consistency.
Figure 6: The average number of predictions made by different rule miners. These measurements are plotted over increasing lattice sizes, and are consistent with the exponential worst-case time complexity of our rule miners. Moreover, they establish $|\mathbf{s}|=10$ as a tractable upper bound for use of our rule miners with modern LLMs.
5.3 Scalable Red Teaming Case Study
We now present a case study focused on a large-scale, real-world scenario in the domain of AI safety. This case study follows an AI safety researcher whose team is developing advanced red teaming methods to safeguard a proprietary RAG LLM system. These methods probe RAG LLMs adversarially in an attempt to identify vulnerabilities, biases, and other harmful behaviors. The user is investigating a vulnerability recently uncovered in the proprietary LLM: malicious instructions can be injected into the RAG prompt through its retrieved sources. Given that users of the proprietary system are able to supply their own RAG sources, patching this vulnerability is imperative to ensure that the system’s safety training is not circumvented at inference time.
The user develops a novel solution that they believe will catch prompt injections and refuse to respond. To test the robustness of this solution, the team applies their AI safety expertise to craft 50 adversarial RAG sources with diverse injection strategies, all attempting to induce a response containing a specific “exploit hash”. Though the user can quickly test each adversarial source independently, they wish to test whether specific combinations of the adversarial sources are effective in exploiting the vulnerability. To efficiently identify any effective combinations, the user seeks rules that associate specific combinations with consistent, successful exploitation.
To generate rules, the user defines a simple output predicate function that checks the output for the exploit hash. The user then provides all 50 adversarial sources to Mono Rule Miner to find all retention rules, but encounters a problem. Given the large number of sources it is searching over, the rule miner is proceeding at a very slow rate and will never terminate. Thinking of ways to reduce the dimensions of the problem, the user realizes that since RAG ultimately concatenates all sources into a single prompt, the 50 sources can also be arbitrarily concatenated or grouped to reduce their count. Once rules are mined for coarse-granularity grouped sources, those not implicated by a rule can be discarded, and finer granularity rules can then be mined by dividing the remaining grouped sources again.
The user writes a script to automate these steps in a divide-and-conquer fashion. Beginning with 50 sources in one grouped source, Mono Rule Miner confirms that the exploit is induced (using 1 LLM inference call). The grouping is therefore divided into 2 grouped sources of size 25 and passed again to Mono, which finds a minimal rule implicating both grouped sources jointly (maximum of $2^{2}$ calls). Since both grouped sources contain one or more sources necessary to induce the exploit, the script continues by dividing each into two grouped sources of size 12 and 13, forming a new set of 4 grouped sources. Mono finds a single rule implicating the first and fourth grouped sources jointly (maximum of $2^{4}$ calls), allowing for the second and third to be discarded. The first and fourth are divided in two again and the algorithm continues similarly; each iteration runs Mono Rule Miner, discards grouped sources not implicated in minimal rules, then divides remaining grouped sources and recurses to solve them.
Once all grouped sources become indivisible (i.e., the original granularity is recovered), the algorithm terminates and returns said sources. After no more than 78 total LLM inference calls, the script provides the user with a single minimal rule at the original source-level granularity: retaining adversarial sources $\left\{\text{D8},\text{D44}\right\}$ consistently induces the exploit. This suggests that D8 and D44 are jointly but not independently effective in bypassing the user’s novel safety solution, perhaps due to interaction effects. With these new insights, the user has been made aware of the exploitative power of these sources when combined, and can study them to refine their solution.
5.4 Quantitative Evaluation: Rule Robustness
Building upon the real-world quantitative evaluation presented in Section 5.2, we present additional quantitative experiments that evaluate the robustness and quality of rules found by our rule miners. We first study the robustness of different “LLMs-under-test”, measuring the precision and recall of our rule miners in a real-world HotpotQA setting. Next, we study the robustness of the “LLM-as-a-judge” output predicate introduced in Section 5.2.1, using expert validation to estimate its precision and recall.
5.4.1 Experimental Setup
In this robustness evaluation, we follow and build directly upon the efficiency evaluation setup described in Section 5.2.1. The goal of the present evaluation is to measure the robustness of our rule miners (which mine “exact” rules) when used with real-world LLMs (whose non-determinism can induce answer stochasticity). We aim to measure and analyze the consistency of the “LLM-under-test” and “LLM-as-a-judge” components, since their stochasticity and propensity for error can lead to false positives and false negatives in their respective functions. Naturally, this evaluation focuses exclusively on the real-world setting, as the synthetic setting (introduced also in Section 5.2.1) deliberately controls for the aforementioned components. Therefore, we once again adopt HotpotQA to conduct a similar multi-hop QA experiment.
Building upon the previous HotpotQA experiment, we adopt the same definition for retention and omission rules, as well as the same implementation for output predicates $O_{consistent}$ and $O_{inconsistent}$. However, the shift in experimental focus from efficiency to robustness necessitates a different parameterization. We propose that the robustness of the LLM components can ultimately be quantified in terms of precision and recall by comparing actual vs. expected rules. After running a rule miner to obtain the “actual” rules, precision measures how many actual were expected (valid), while recall measures how many expected rules were actually output. We construct ground truth “expected” rules from the HotpotQA QA pairs to facilitate this computation; retention rules must retain both necessary sources to induce a consistent answer, while omission rules must omit one or more necessary sources to induce an inconsistent answer.
Using these metrics, we proceed similarly to the steps described in Section 5.2.1. We draw 50 QA pairs from HotpotQA training set at random, and fix each with exactly 5 sources: both necessary sources and three additional distractor sources. We pose these questions to three different commercial RAG LLMs, each initialized with a temperature of 1.0: OpenAI’s GPT-5 Mini (gpt-5-mini-2025-08-07), Google’s Gemini 2.5 Flash (gemini-2.5-flash), and Anthropic’s Claude Haiku 4.5 (claude-haiku-4-5-20251001). This enables us to contrast robustness across several commercial LLMs of similar capability.
We task Dual Rule Miner with finding all rules (of both types) for each LLM posed with each question, but do so repeatedly to obtain 10 sampled LLM answers for each. Due to their inherent stochasticity, LLMs may produce any number of different outputs when posed repeatedly with the same input prompt, potentially contradicting itself. Many works in both the machine learning and LLM literature have formalized inference strategies that embrace this reality (e.g., ensemble methods [ensemble_methods, llm_ensemble]), often improving overall accuracy by sampling multiple answers and computing a single final answer (e.g., via majority voting). Building upon these works, we compute a majority answer for each sample size in the range of 1 to 10, assessing the sample group as consistent only when 50% or more individual samples are judged consistent (see Appendix A for more detail). We then contrast and analyze the evolution of robustness over these increasing sample sizes, thereby measuring the robustness of our “exact” rules in the face of answer stochasticity.
5.4.2 Experimental Results: “LLM-under-test” Robustness
In Figure 7, we present the robustness of the different “LLMs-under-test”, reporting on four pairings: precision and recall over both retention and omission rules. In general, the chosen LLMs demonstrate similar robustness within each pairing, and across all samples sizes therein. With the exception of omission rule precision, for which all LLMs score approximately 100%, the results indicate that Gemini 2.5 Flash is the highest performing, followed closely by Claude Haiku 4.5 and then GPT-5 Mini. Despite using a moderate LLM temperature of 1.0 to permit stochasticity, all robustness measurements remain nearly constant as more samples are considered for the majority vote. This indicates that, despite the possibility of contradicting each other, the stochastic outputs of LLMs can (at least in similar QA settings) remain sufficiently consistent to produce exact rules, regardless of sample size.
In Figures 7(a) and 7(b), we observe that omission rules exhibit near-perfect precision but significantly lower recall. This behavior reflects the fact that omission rules are easier to satisfy; in theory, omitting a single necessary source will suffice to induce an inconsistent answer. Therefore, more omission rules should exist than retention rules, as established in Section 5.2.2. It follows that the omission rules found are almost always valid (high precision), since many exist. Consequently, there are even more omission rules that may not hold under a fallible LLM, resulting in some being missed (lower precision) in our “exact rule” formulation.
(a) Precision of omission rules generated.
(b) Recall of omission rules generated.
(c) Precision of retention rules generated.
(d) Recall of retention rules generated.
Figure 7: The precision and recall of rules generated using different “LLMs-under-test” over an increasing number of answer samples.
In Figures 7(c) and 7(d), we observe the opposite trend: retention rules exhibit near-perfect recall but significantly lower precision. Although all necessary sources should be required to derive the ground truth answer, we observe many instances where predicted answers are consistent nonetheless. We contend that this exceptional behavior stems from data quality issues in the HotpotQA dataset, as well as the LLMs’ unexpected propensity to leverage their pre-trained knowledge and “guess”. Consequently, many source combinations satisfy the retention rule criteria in practice despite omitting one or both of the necessary sources (low precision). Likewise, the presence of all necessary sources almost always induces a consistent answer, leading to the discovery of most expected retention rules (high recall).
Diving deeper to rationalize this unexpected behavior, we note that several flaws have been documented in HotpotQA questions, such as sources being mislabeled as non-relevant [ni-etal-2021-mitigating, thakur-etal-2025-hard] and the inadvertent introduction of reasoning shortcuts [jiang-bansal-2019-avoiding, chen-durrett-2019-understanding]. Qualitatively analyzing our experimental results, we find evidence that these flaws contributed to the validation of unexpected retention rules (false positives) and invalidation of expected omission rules (false negatives).
Specifically, we find that some distractor sources inadvertently contain information relevant to the question, and can be used to induce a consistent final answer without all necessary sources. For example, one question asks “Which rock group was formed earlier, Beastie Boys or The Shins?” There are two necessary sources, one for each band; each states the year the group was formed (1981 and 1996 respectively). However, some distractors are unexpectedly relevant by virtue of establishing pre-1996 release dates for some Beastie Boys songs. So long as “The Shins” necessary source is retained, a consistent answer can be inferred by also retaining any of these “release date” distractors.
Furthermore, we note that general HotpotQA data quality issues and inconsistencies may also affect our results. For instance, one question used in our experiments asks “When was the composer who wrote the music for Euryanthe born?”, but the ground truth answer “31 January 179719” is malformed. In this same example, the first necessary source establishes that “Carl Maria von Weber” wrote the music for “Euryanthe”, but the second necessary source states only the malformed birth date of the composer “Franz Peter Schubert”, who did not write “Euryanthe”. To illustrate another issue, we cite the question “Which profession does Boris Barnet have in common with Baltasar Kormákur?”, whose ground truth answer is “film director”. Despite the fact that HotpotQA questions should have a unique answer, we find that the two necessary sources indicate “actor” as another valid answer.
In cases where necessary information was truly absent, the LLM often chose to defy its instructions and “guess” the answer instead of strictly inferring it. Many of these guesses were consistent (and judged as such), perhaps due to the fact that the commercial LLMs tested likely trained on the same Wikipedia data from which HotpotQA was constructed. The remaining inconsistent guesses accounted for much of the stochasticity observed across answer samples. We note that answer diversity can be expected to increase as temperature is increased, and decrease as temperature is decreased. However, since diversity is offset by our “majority answer” sampling, we would not expect rule validity to be significantly impacted if our experiments were repeated with a temperature of 0.
5.4.3 Experimental Results: “LLM-as-a-judge” Robustness
To assess the reliability of the “LLM-as-a-judge” used across all of our quantitative experiments, we conduct an evaluation of its precision and recall utilizing ground truth answers. We randomly sample 100 judgments made by the “LLM-as-a-judge” during the robustness experiments (inspecting all 10 answer samples), drawn randomly across all QA pairs and “LLMs-under-test”. Each judgment is reviewed for correctness by an expert validator, who assesses its correctness by comparing the consistency of the predicted answer w.r.t. the HotpotQA ground truth answer. By tracking the number of examples that were accurately and inaccurately judged as “consistent” and “inconsistent”, we can compute the precision and recall of the “LLM-as-a-judge”.
Across the evaluated judgments, the “LLM-as-a-judge” achieves precision of 98.89% and recall of 100%, indicating that nearly all answers judged “consistent” were correct, and all truly “consistent” answers were correctly judged as such. The false positive rate is 0.90%, indicating that extremely few “inconsistent” answers were judged “consistent” by mistake. These results indicate that the judge is highly reliable at identifying answers consistent with the ground truth answer, and makes few mistakes in this QA setting. Moreover, we observe that the “LLM-as-a-judge” judges consistency at a semantic level, affording a level of nuance that makes it robust to unexpected inconsistencies.
Specifically, we observe that the “LLM-as-a-judge” is robust in identifying alternative answers that are consistent with the ground truth answer provided. For example, the HotpotQA question “Kevin Olekaibe plays for the team that competes in what league?” is labeled with the ground truth answer “NBA G League”. This league is known more generally as the NBA’s developmental league, and has been renamed multiple times (most recently to “NBA G League”, named after its current sponsor “Gatorade”). This question introduces ambiguity, as previous names for the league may or may not be consistent with ground truth. Expert validation finds that the “LLM-as-a-judge” exhibits semantic understanding of the answers in the context of the question, and judges accordingly. In one response to this question, the “LLM-as-a-judge” correctly inferred that “NBA Development League” is a more general term to describe the “NBA G League”, and correctly judged it as “consistent”.
Beyond robustness to diverse answers, we also observe robustness to diverse LLM responses. Our experiments show that Claude Haiku 4.5, despite being limited to 100 output tokens and prompted to output only the final answer, frequently outputs reasoning text that exhausts the allotted output tokens and is eventually truncated (i.e., before establishing a final answer). In most cases, these responses should be judged as inconsistent since a final answer was not clearly established, but reasoning sufficiently conclusive and aligned with ground truth could qualify as consistent. We note that across truncated responses observed via expert validation, those misjudged as “consistent” typically proposed the ground truth answer as a candidate answer, but were truncated shortly before articulating a final answer. These situations demonstrate how judgment of reasoning-heavy or partial answers is nuanced and potentially subjective, but the “LLM-as-a-judge” excels in this setting. Expert validation confirms that these truncated responses are rarely misjudged, as evidenced by the low false positive rate.
6 Conclusions
We proposed the first formulation of rule-based explanations for retrieval-augmented LLM systems. We presented two complementary variants that leverage the concept of feature ablation to link certain input features to the satisfaction of arbitrary output predicates. We implemented efficient algorithms for mining these rules, which make use of pruning and early termination. By way of quantitative and qualitative evaluations, we demonstrated the efficiency of our algorithms and the utility of our explanations, as applied in real-world RAG scenarios.
Our novel formulations and algorithms invite promising directions for future work. For instance, a relaxed version of our formulation could model rules that hold approximately (i.e., with some minimum confidence) rather than exactly (i.e., with certainty), which could help model the nuances of LLM non-determinism. Furthermore, with our rule mining algorithms, HotpotQA and similar benchmarks present an opportunity for novel benchmarking studies. Since expected rules can be derived from HotpotQA’s ground truth answers and supporting sources, a study could empirically compare LLMs on the basis of whether rules expected to hold actually do. Similarly, this ground truth data could be used to assess how often an LLM hallucinates answers not present in retrieved sources, or how often it deviates from RAG prompt instructions.
Declarations
6.1 Ethics Approval and Consent to Participate
Not applicable.
6.2 Consent for Publication
Not applicable.
6.3 Availability of Data and Material
Implementations of our proposed algorithms, along with all code and data used to conduct our quantitative experiments, are available in a public repository hosted at https://github.com/joelrorseth/rag-llm-rule-miner. The synthetic data used in our quantitative experiments is generated by the experiments at runtime in a fully reproducible manner. The HotpotQA dataset [hotpotqa] used in our quantitative experiments has been made available publicly by its authors, but instructions for downloading it are included in our repository.
6.4 Competing Interests
The authors declare no competing interests.
6.5 Funding
This research was supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC) through a PGS D scholarship (awarded to the first author) and Discovery Grant funding (awarded to multiple co-authors). One author was supported by the Connected Minds program, funded by the Canada First Research Excellence Fund (CFREF).
6.6 Author Contributions
Conceptualization: Joel Rorseth, Parke Godfrey, Lukasz Golab, Divesh Srivastava and Jarek Szlichta. Methodology, software, data curation, formal analysis, investigation, visualization, and writing – original draft: Joel Rorseth. Supervision and writing – review and editing: Parke Godfrey, Lukasz Golab, Divesh Srivastava and Jarek Szlichta. All authors read and approved the final manuscript.
Appendix A LLM Prompt Templates for Experimental Evaluation
Each of the “LLMs-under-test” in the quantitative evaluations is posed with a prompt constructed from the template in Listing 1. The text for the given question and knowledge sources is inserted into the corresponding template variables, along with additional newlines to delineate individual sources. If no sources are given, prediction is avoided any “N/A” is returned. For all experiments and LLMs, the temperature parameter is set to 1.0, reasoning effort to the minimum allowed value, and maximum number of output tokens to 100. Before drawing HotpotQA questions for the experiments, we discard those whose ground truth answer has more than 100 characters, ensuring that the output token limit can be respected.
Listing 1: “LLM-under-test” Prompt Template
You are a question answering agent that uses knowledge sources to infer answers to questions.
You will be given a question and a set of knowledge sources.
Your goal is to determine the correct answer to the question.
**Question**
{question}
**Sources**
{sources}
**Task**
Determine the correct answer to the question using the knowledge sources.
If an answer cannot be inferred from the knowledge sources, return N/A.
**Output**
Return only the final answer, and nothing else.
Each answer generated by an “LLM-under-test” is judged w.r.t. its respective ground truth answer, first using string matching to identify near-exact matches, then using “LLM-as-a-judge” if no match was found. The string matching procedure checks for equality between a candidate and ground truth answer after performing case folding, punctuation stripping, NFKD (Normalization Form Compatibility Decomposition), and ASCII encoding to remove diacritics. If the strings do not match, the candidate and ground truth answer are posed to an LLM (along with the question) to classify their semantic equivalence (“consistency”). These pieces of information are substituted into the prompt template in Listing 2, which is subsequently posed to the LLM.
In the robustness experiments presented in Section 5.4, we pose the same prompt to each “LLM-under-test” up to 10 times. We compute a final judgment for the sampled answers using binary majority aggregation: the “final answer” is judged “consistent” only if 50% or more of the sampled answers are judged “consistent”. Since sample sizes are tested in increasing order, we cache previous responses for all “LLMs-under-test” to improve experimental efficiency, computing only a single new answer sample in each round. To further improve efficiency, all sampled answers (for the same question) are batched into a single “LLM-as-a-judge” prompt and assessed in one shot. The prompt therefore instructs the LLM to assess the equivalency of each candidate (sampled answer) against the ground truth answer.
Listing 2: “LLM-as-a-judge” Prompt Template
You are a question answering grading agent that assesses the equivalence of candidate answers against ground truth answers.
You will be given a question, ground truth that expresses the true answer, and a list of candidate answers.
Your goal is to assess whether each candidate is equivalent to the ground truth answer.
**Question**
{question}
**Ground Truth**
{ground_truth_answer}
**Candidate Answers**
{candidate_answers}
**Task**
Determine whether each candidate answer is equivalent to the ground truth answer, with respect to the given question.
Note that while the ground truth may include additional details, the candidate need only be equivalent to the final answer expressed therein.
**Output**
Return a list of booleans, one for each candidate answer in the same order they were given.
The boolean should be True if the corresponding candidate answer is equivalent to the ground truth answer, and False otherwise.
Appendix B Rule Count for Efficiency Evaluation
(a) Rule count as evaluated on a real LLM tasked with answering HotpotQA questions.
(b) Rule count as evaluated in a synthetic setting that controls for stochasticity and consistency.
Figure 8: The average number of valid rules found by different rule miners. These measurements are plotted over increasing lattice sizes, showing that our Mono and Dual rule miners find more rules as more sources are considered.
In Figure 8, we illustrate the number of rules found by each rule miner in our quantitative efficiency experiment in Section 5.2. The difference in HotpotQA vs. synthetic trends, discussed in Section 5.2, are evident here as well. Under our RAG QA formulation, more omission rules should exist than retention rules (as indicated in Figure 8(b)) since they are easier to satisfy. However, in the practical HotpotQA setting using a commercial LLM, more retention rules are valid than expected and fewer omission rules hold (as indicated in Figure 8(a)). The trends are asymptotically proportional to the scalability plots in Figure 6, which is why they appear in the Appendix.